# Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection

### v2

@@ 3.3 Selecting Top-$k$ Adversarial Prompts
Using all the style-conversion prompt candidates generated by the LLM can be computationally intensive, and not all prompts may be suitable for augmentation. For example, the set of style-conversion prompts used for augmentation should each provide adversarial perturbations that confuse the detector (i.e., adversarialness). Additionally, while altering the textual style, the prompts should not change the meaning of the sentences to prevent label noise (i.e., coherency). Furthermore, the more diverse the set of conversion directions covered by the prompts, the more efficient the augmentation process (i.e., diversity).
@@
$\mathbf{z}$ $=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{d}_{i}\in\mathcal{B}}{g(\mathbf{d}_{i})},\ \ \ \mathbf{z}\textquoteright=\frac{1}{|\mathcal{B}\textquoteright|}\sum_{\mathbf{d}^{c}_{i}\in\mathcal{B}\textquoteright}{g(\mathbf{d}^{c}_{i})}$
$\mathbf{z}_{c}$ $=\mathbf{z}\textquoteright-\mathbf{z}$ (2)
@@
Here, $\mathbf{z}_{c}$ represents the average change in embedding direction due to the conversion prompt $c$. We then calculate the adversarialness scale $s^{c}_{\text{adv}}$ and coherency scale $s^{c}_{\text{coh}}$ for each conversion prompt $c$, and rescale $\mathbf{z}_{c}$ accordingly (i.e., $\mathbf{\hat{z}}_{c}=\mathbf{z}_{c}\times s^{c}_{\text{adv}}\times s^{c}_{\text{coh}}$). Finally, we use the $k$-means++ initialization method Arthur et al. (2007) on these rescaled vectors to select $k$ prompts. The $k$-means++ initialization method helps select a diverse set of prompts that are adversarial and coherent, by choosing samples that are as far apart as possible Ash et al. (2019). The details of each scale are described below.
Adversarialness scale.
@@
To measure how adversarial a given style-conversion prompt is to the detector, we define a new adversarialness scale similar to the confusion score defined in the previous section. Given the converted batch $\mathcal{B}^{c}$ by conversion prompt $c$, the adversarialness scale $s^{c}_{\text{adv}}$ is defined as $-1.8\cdot|\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})-0.5|+1$. This value increases as the AUC approaches 0.5, indicating that the prediction is more random. The coefficient 1.8 ensures that the scale ranges between 0.1 and 1, preventing it from being zero.
Coherency scale.
@@
To verify that the converted text retains the same content as the original text, we check the similarity in meaning between the text pairs using an LLM. Specifically, we create sample pairs from $\mathcal{B}$ and the converted subset $\mathcal{B}^{c}$, and inquire the LLM about the percentage of pairs that it considers to have the same meaning. This percentage is used as the coherency scale $s^{c}_{\text{coh}}$. Like the adversarialness scale, this value is rescaled to range between 0.1 and 1.
@@
Finally, the selected style-conversion prompts are applied to the input texts of the entire dataset $\mathcal{D}$ to create augmented samples. These augmented samples are then used alongside the original samples to train the detector $f$. The detector is trained using binary cross-entropy loss. These prompt generation and selection processes are repeated over multiple training rounds.
4 Experiment
@@
Table 1: Performance comparison with AdStyle in two different scenarios—Attack, where style-conversion attacks are performed, and Clean, where no attack is performed—across three fake news datasets. For the attack scenario, we report the average AUC of four style-conversion attacks. The best results are marked in bold. Our model demonstrates a significant performance improvement over all text-based fake news detectors in both style-conversion attack and clean scenarios.
@@
Table 2: Performance comparison of ablations on the Politifact dataset. The results show the impact of style-conversion attacks using four different publishers (i.e., CNN, The New York Times, The Sun, and National Enquirer) and a clean scenario (i.e., Clean) where no attack is performed. Any modification or removal of model components leads to decreased performance.
@@
Adv. only Selection 0.9035 0.9343 0.9318 0.9288 0.9405
w/o Uncertainty 0.9039 0.9333 0.9320 0.9402 0.9473
w/o Similarity 0.9193 0.9315 0.9297 0.9430 0.9409
@@
We evaluate the robustness of AdStyle under diverse style-conversion attacks across multiple datasets, comparing it with contemporary baselines. Then, we analyze the impact of model components on overall performance. Lastly, we conducted a qualitative analysis to investigate the characteristics of the style-conversion prompts generated by the LLM. Due to limited space, additional analysis, such as the performance evaluation over wider variety of paraphrasing attacks and comparisons with LLM-based zero-shot and in-context learning baselines, has been included in the Appendix.
4.1 Performance Evaluation
Dataset.
Our experiments use three real-world fake news benchmark datasets. We utilize PolitiFact and Gossipcop, drawn from the FakeNewsNet benchmark Shu et al. (2020), which focus on political claims and celebrity rumors, respectively. Additionally, we incorporate Constraint Felber (2021), a dataset specifically addressing COVID-19 related social media posts. Each dataset is randomly split into an 80% training set and a 20% test set. Detailed dataset statistics are provided in Table 3 of the Appendix.
Attack settings.
To assess robustness against style conversion attacks, we employ LLM-empowered techniques to reframe the test set using a variety of style conversion prompts, as illustrated in Figure 3. Following the original literature Wu and Hooi (2023), we use four well-known daily news sources as [publisher name]: CNN, The New York Times, The Sun, and National Enquirer. CNN and The New York Times are recognized for their reputable journalism, while The Sun and National Enquirer are characterized by their tabloid style. We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences, with the temperature set to 0 and the top-p value set to 1 by default. We conduct experiments using 10%, 25%, and 100% of the complete dataset to observe the effect of augmentation across different dataset sizes for training. AUC is used as our evaluation metric.
Figure 3: Prompt for style conversion. The “publisher name" part will be filled with the name of a representative publisher (e.g., newspaper or journal), and the “news article" part will contain the original news text.
Baselines.
@@ 4.1 Performance Evaluation
We have implemented several existing text-based fake news detection strategies as baselines: (1) Vanilla, a text-based fake news detector using conventional binary cross entropy objective; (2) UDA, introduces consistency regularization objective between original text and diverse augmented variations. Xie et al. (2020); (3) RADAR, utilizes an adversarially trained paraphraser to generate augmented version of input sentences Hu et al. (2023); (4) ENDEF, mitigates entity bias in fake news data through causal learning Zhu et al. (2022); (5) SheepDog, introduces predefined style-conversion prompts to augment the styles of input text via LLM Wu and Hooi (2023). We follow the original paper’s setting and details for baseline implementations.
Implementation details.
@@
All models are evaluated under uniform experimental conditions to ensure fair comparison. This consistency extends to the choice of backbone network, optimizer, and learning rate. We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences and measuring coherency, with the temperature set to 0 and the top-p value set to 1 by default. Our training process for the detector is conducted over 10 rounds, with one training epoch per round. When evaluating the style-conversion prompts, we randomly selected 30 samples from the training dataset to apply augmentation (i.e., $M=30$). In each round, 30 prompt candidates were generated by the LLM (i.e., $S=30$), from which 3 were chosen for augmentation (i.e., $k=3$) by our selection strategy. The training utilizes the AdamW optimizer with a learning rate of 1e-5 and a batch size of 8. For measuring diversity, the text embeddings are generated using a pre-trained BERT-based uncased model from HuggingFace Transformers. Two V100 GPUs were utilized for all experiments.
Result.
Table 1 compares the performance of detector algorithms in both clean scenarios, where no attack is performed, and adversarial scenarios, where style-conversion attacks are applied. Due to space limitations, the average AUC results under the four different style-conversion attacks are reported. Detailed results for each prompt can be found in the Appendix (See Table 6 in the Appendix). According to the results, AdStyle consistently outperforms the baselines across all cases, including both clean and adversarial scenarios. This demonstrates that our augmentation strategy enhances both the robustness and generalizability of the detector. Notably, our model’s effectiveness is more pronounced when compared to other baselines, especially with smaller datasets. Figure 4 shows the AUC for each style-conversion attack over different rounds. The performance gradually improves as the rounds progress, indicating that continually discovering adversarial augmentations is beneficial.
@@
Figure 4: Performance changes across rounds on the PolitiFact dataset for four style-conversion attacks. The x-axis represents the training rounds, and the y-axis represents the detector’s AUC. For all attacks, the detector’s performance improved as the rounds progressed.
4.2 In-Depth Performance Analysis
We have demonstrated AdStyle’s effectiveness. In this section, we further examine the contribution of each component and provide a qualitative analysis on our adversarial style-conversion prompts.
Ablation study.
@@ 4.2 In-Depth Performance Analysis
The proposed method integrates two main modules: adversarial style-conversion prompts generation and selection. To evaluate their individual contributions, we conduct experiments where we either remove each component or substituted it with an alternative within the full model. This results in six distinct configurations for analysis: (1) Full Components: Our complete method with all components; (2) Random Selection: The method that randomly select prompts from candidates instead of using our selection strategy; (3) Adversarial only Selection: The method that selects top-$k$ adversarial prompts, not considering diversity and coherence; (4) w/o Adversarialness: The method that omits the adversarialness score in our selection strategy. (5) w/o Coherency: The method that omits the similarity score in our selection strategy. (6) w/o Score trajectory: The method without score trajectory component for the style-conversion prompt generation.
Table 2 demonstrates that omitting any component leads to a decrease in performance for certain style conversions attack. While omitted due to space limitation, it can be confirmed that our sampling strategy converges faster than other sampling strategies (See Figure 8 in Appendix).
(a) Diversity
(b) Adversarialness
(c) Coherency
Figure 5: Qualitative analyses of AdStyle. (a) Histogram of diversity in embedding ($\mathbf{z}_{c}$) of conversion prompts sampled by our method and adversarial only selection. (b) Histogram of adversarialness scores ($s_{adv}^{c}$) for selected and unselected prompts (c) Histogram of coherency scores for selected and unselected prompts
Qualitative analyses.
First, we verified that our sampling strategy effectively selects diverse style-conversion prompts with high adversarialness and coherency. Figure 5(a) visualizes the diversity of prompts selected by our method compared to those chosen solely based on adversarialness scores (i.e., Adversarial-only Selection, the third model in our ablation study). We measure diversity using the average cosine similarity of every pair of selected prompts’ embeddings, $\mathbf{z}_{c}$ (Eq. 2):
$\text{Diveristy}(\mathcal{C})$ $=1-{1\over|\mathcal{C}|}\sum_{(c_{i},c_{j})\in\mathcal{C}}{\text{sim}(\mathbf{z}_{c_{i}},\mathbf{z}_{c_{j}})},$ (3)
Figure 5(b) and 5(c) illustrate the Adversarialness and Coherency, respectively, of our selected prompts compared to remaining unselected prompts. Adversarialness was measured using the $s_{adv}^{c}$ score (Section 3.3), and coherency was calculated as the cosine similarity between the original text and its augmented version using semantic BERT embeddings Chanchani and Huang (2023). We can also observe that, for both metrics, prompts selected through AdStyle exhibit higher values compared to unselected prompts. This suggests that our strategy effectively selects prompts by considering both adversarialness and coherency.
Next, we qualitatively analyzed the characteristics of our style-conversion prompts. Figure 6 shows examples of prompts generated and selected by AdStyle. Previous work relied on prompts that were more conventional, such as “neutral" or “sensational." However, our prompts included creative phrases beyond typical human suggestions, which proved to be more adversarial to the detector. Currently, we use a fixed format for the initial set of prompts, which results in the generation of prompts with a similar format. We expect that using a more diverse initial set of prompts will allow the LLM to optimize prompts within a wider search space.
Round 0 P1: Rewrite the following article in a nonsensical and absurdly exaggerated tone with a hint of horror P2: Rewrite the following article in a sarcastic and mocking tone P3: Rewrite the following article in a chaotic and disorganized tone Round 1 P1: Rewrite the following article in a haunting and macabre tone with a sense of impending horror and madness P2: Rewrite the following article in a cryptic and enigmatic tone P3: Rewrite the following article in a malevolent and apocalyptic tone with a sense of impending doom and destruction, while also incorporating elements of surrealism and existential dread
Figure 6: Examples of our adversarial style-conversion prompts selected for the PolitiFact dataset
5 Conclusion
This paper presents a robust fake news detection method that withstands various style-conversion and paraphrasing attacks through adversarial style-conversion. Unlike traditional detectors that use predefined, agnostic augmentations, AdStyle employs tailored augmentations that shift samples in the direction of the detector’s current decision boundary using style-conversion prompts, functioning similarly to adversarial noise. Among the various prompt candidates generated by the LLM, we selected an efficient set of prompts for training by considering diversity, coherency, and adversarialness. As a result, we were able to train a detector that exhibits high robustness and generalizability against a wide range of attacks. We believe that our work helps filter fake news and contribute to a better exchange of information in online.
Appendix A APPENDIX
A.1 Data & Implementation Details
Dataset overview.
This paper utilizes a diverse range of datasets sourced from publicly available fake news detection resources, selected for their unique topical focuses. We utilize PolitiFact and Gossipcop, drawn from the FakeNewsNet benchmark Shu et al. (2020), addressing political claims and celebrity rumors respectively. Furthermore, we include Constraint Felber (2021), a dataset focused on COVID-19 related social media posts. Each dataset is randomly divided into an 80% training set and a 20% test set. Detailed statistics for these datasets can be found in Table 3.
Table 3: Statistics of fake news datasets.
Dataset PolitiFact GossipCop Constraint
# of News Articles 774 7,916 8,418
# of Real News 399 3,958 4,406
# of Fake News 375 3,958 4,012
Implementation details.
For consistency, experiments replicating existing baselines maintained fixed settings for learning, backbone network, and other relevant parameters. Augmented variations for UDA were generated using back-translation through German. RADAR employed adversarial training to learn paraphrases, utilizing the maximization of the binary cross-entropy loss of a fake news detector as a reward signal, with the T5-large model serving as the paraphraser. Named entities for ENDEF were extracted using the bert-base-NER model from Hugging Face. For SheepDog, augmentations were generated following the original paper’s methodology, using the prompt format illustrated in Figure 7 and four tones: "objective and professional," "neutral," "emotionally triggering," and "sensational". We also utilized the same four tones and prompt format as our predefined initial prompt set. Two V100 GPUs were utilized for all experiments.
Rewrite the following article in a/an [tone]: [news article]
Figure 7: Prompt for style-conversion. The “tone" part will be filled with the desired tone, and the “news article" part will contain the original news text.
A.2 Further Results on Evaluation.
Result on selection strategies across rounds.
Figure 8 shows the AUC for each selection strategy over different training rounds. Random Selection leads to slow convergence and unstable training, while Adversarial-only Selection, limited by its narrow focus on adversarial prompts, can negatively impact performance in later stages of training due to a lack of diversity in the generated prompts. Notably, without score trajectories, the model struggles to generate prompts that contribute effectively to the training process.
@@ A.1 Data & Implementation Details
Figure 8: Performance of selection strategies over training rounds on the PolitiFact dataset. The x-axis represents the training rounds, while the y-axis represents the detector’s AUC.

### v3

@@ 3.3. Selecting Top-$k$ Adversarial Prompts
Using all the style-conversion prompt candidates generated by the LLM can be computationally intensive, and not all prompts may be suitable for augmentation. For example, the set of style-conversion prompts used for augmentation should each provide adversarial perturbations that confuse the detector (i.e., adversarialness). While altering the textual style, the prompts should not change the meaning of the sentences to prevent label noise (i.e., coherency). The more diverse the set of conversion directions covered by the prompts, the more efficient the augmentation process (i.e., diversity).
@@
$\mathbf{z}$ $=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{d}_{i}\in\mathcal{B}}{g(\mathbf{d}_{i})},\ \ \ \mathbf{z}’=\frac{1}{|\mathcal{B}’|}\sum_{\mathbf{d}^{c}_{i}\in\mathcal{B}’}{g(\mathbf{d}^{c}_{i})}$
(Eq. 2) $\mathbf{z}_{c}$ $=\mathbf{z}’-\mathbf{z}$
@@
Here, $\mathbf{z}_{c}$ represents the average change in embedding direction due to the conversion prompt $c$. We then calculate the adversarialness scale $s^{c}_{\text{adv}}$ and coherency scale $s^{c}_{\text{coh}}$ for each conversion prompt $c$, and rescale $\mathbf{z}_{c}$ accordingly (i.e., $\mathbf{\hat{z}}_{c}=\mathbf{z}_{c}\times s^{c}_{\text{adv}}\times s^{c}_{\text{coh}}$). Finally, we use the $k$-means++ initialization method (Arthur et al., 2007) on these rescaled vectors to select $k$ prompts. The $k$-means++ initialization method helps select a diverse set of prompts that are adversarial and coherent, by choosing samples that are as far apart as possible (Ash et al., 2019). The details of each scale are described below.
@@
Adversarialness scale ($s^{c}_{\text{adv}}$). To measure how adversarial a given style-conversion prompt is to the detector, we newly define the adversarialness scale similar to the confusion score defined in the previous section. Given the converted batch $\mathcal{B}^{c}$ by conversion prompt $c$, the adversarialness scale $s^{c}_{\text{adv}}$ is defined as:
(Eq. 3) $s^{c}_{\text{adv}}=-1.8\cdot|\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})-0.5|+1.$
This value increases as the AUC approaches 0.5, indicating that the prediction is more random. The coefficient 1.8 ensures that the scale ranges between 0.1 and 1, preventing it from being zero.
@@
Coherency scale ($s^{c}_{\text{coh}}$). To verify that the converted text retains the same content as the original text, we check the similarity in meaning between the text pairs using an LLM. We create sample pairs from $\mathcal{B}$ and the converted subset $\mathcal{B}^{c}$, and inquire the LLM about the percentage of pairs that it considers to have the same meaning. This percentage is used as the coherency scale $s^{c}_{\text{coh}}$. Like the adversarialness scale, this value is rescaled to range between 0.1 and 1.
@@
Finally, the selected style-conversion prompts via our score and $k$-means++ initialization method are applied to the input texts of the entire dataset $\mathcal{D}$ to create augmented samples. These augmented samples are then used alongside the original samples to train the detector $f$. The detector is trained using binary cross-entropy loss. These prompt generation and selection processes are repeated over multiple training rounds.
4. Experiment
We evaluate the robustness of AdStyle under diverse style-conversion attacks across multiple datasets, comparing it with contemporary baselines. We then analyze the impact of various model components on overall performance. A comprehensive evaluation is also performed on a broader range of paraphrasing attacks, including comparisons with LLM-based zero-shot and in-context learning baselines. Lastly, a qualitative analysis is conducted to explore the characteristics of the generated style-conversion prompts.
4.1. Performance Evaluation
Dataset. We use three real-world fake news benchmark datasets: (1) PolitiFact and (2) Gossipcop, released by FakeNewsNet benchmark (Shu et al., 2020), which focus on political claims and celebrity rumors, respectively, and (3) Constraint (Felber, 2021), a dataset of social media posts on COVID-19. Datasets were split into an 80% training set and a 20% test set, as described in Table 1.
Table 1. Statistics of fake news datasets.
Dataset PolitiFact GossipCop Constraint
# of News Articles 774 7,916 8,418
# of Real News 399 3,958 4,406
# of Fake News 375 3,958 4,012
@@ 4.1. Performance Evaluation
Table 2. Performance comparison with AdStyle in two different scenarios—Attack, where style-conversion attacks are performed, and Clean, where no attack is performed—across three fake news datasets. For the attack scenario, we report the average AUC of four style-conversion attacks. The best results are marked in bold. The values 0.1, 0.25, and 1 indicate the proportion of the dataset used for training relative to the full dataset. Our model demonstrates a significant performance improvement over all text-based fake news detectors in both style-conversion attack and clean scenarios.
@@
Attack settings. To assess robustness against style conversion attacks, we employ LLM-empowered techniques to reframe the test set using a variety of style conversion prompts, as illustrated in Figure 3. Following the original literature (Wu and Hooi, 2023), we use four prominent daily news sources as [publisher name]: CNN, The New York Times, The Sun, and National Enquirer. CNN and The New York Times were chosen as representative news outlets recognized for their reputable journalism, while The Sun and National Enquirer are characterized by their tabloid style. We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences, with the temperature set to 0 and the top-p value set to 1 by default. Experiments are conducted using 10%, 25%, and 100% of the complete dataset to observe the impact of augmentation on training performance, measured by AUC, across various dataset sizes.
Figure 3. Prompt for style conversion. The “publisher name” part will be filled with the name of a representative publisher (e.g., newspaper or journal), and the “news article” part will contain the original news text.
Table 3. Performance comparison of ablations on the Politifact dataset. The results show the impact of style-conversion attacks using four different publishers (i.e., CNN, The New York Times, The Sun, and National Enquirer) and a clean scenario (i.e., Clean) where no attack is performed. Any modification or removal of model components leads to decreased performance.
@@
Class Prompt 0.9131 0.9272 0.9333 0.9313 0.9471
Adversarial only Selection 0.9035 0.9343 0.9318 0.9288 0.9405
w/o Adversarialness 0.9039 0.9333 0.9320 0.9402 0.9473
w/o Coherency 0.9193 0.9315 0.9297 0.9430 0.9409
@@
Baselines. We implemented several existing text-based fake news detection strategies as baselines: (1) Vanilla, a text-based fake news detector using conventional binary cross entropy objective; (2) UDA, introduces consistency regularization objective between original text and diverse augmented variations. (Xie et al., 2020); (3) RADAR, utilizes an adversarially trained paraphraser to generate augmented version of input sentences (Hu et al., 2023); (4) ENDEF, mitigates entity bias in fake news data through causal learning (Zhu et al., 2022); (5) SheepDog, introduces predefined style-conversion prompts to augment the styles of input text via LLM (Wu and Hooi, 2023). We follow the original paper’s setting and details for baseline implementations.
@@
Implementation details. All models are evaluated under uniform experimental conditions to ensure fair comparison. This consistency extends to the choice of backbone network, optimizer, and learning rate. We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences and measuring coherency, with the temperature set to 0 and the top-p value set to 1 by default. The training process for the detector is conducted over 10 rounds, with one training epoch per round. When evaluating the style-conversion prompts, we randomly selected 30 samples from the training dataset to apply augmentation (i.e., $M=30$). In each round, 30 prompt candidates were generated by the LLM (i.e., $S=30$), from which 3 were chosen for augmentation (i.e., $k=3$) by our selection strategy. The training utilizes the AdamW optimizer with a learning rate of 1e-5 and a batch size of 8. For measuring diversity, the text embeddings are generated using a pre-trained BERT-based uncased model from HuggingFace Transformers. Two V100 GPUs were utilized for all experiments.
@@
Figure 4. Performance changes across rounds on the PolitiFact dataset for four different style-conversion attacks. The x-axis represents the training rounds, and the y-axis represents the detector’s AUC. For all attacks, the detector’s performance improved as the rounds progressed.
Result. Table 2 shows results comparing the performance of detector algorithms in both clean scenarios, where no attack is performed, and adversarial scenarios, where style-conversion attacks are applied. Due to space limitations, the average AUC results under the four different style-conversion attacks are reported. We find that AdStyle consistently outperforms the baselines, including both clean and adversarial scenarios. This demonstrates that our augmentation strategy enhances both the robustness and generalizability of the detector. The model is particularly effective when compared to other baselines and for smaller datasets. Figure 4 shows the AUC for each style-conversion attack over different rounds. The performance gradually improves as the rounds progress, indicating that continually discovering adversarial augmentations is beneficial.
4.2. Component Analysis
@@
We now examine the contribution of each component on our adversarial style-conversion prompts. The proposed method integrates two main modules: adversarial style-conversion prompts generation and selection. To evaluate their individual contributions, we conduct experiments where we either remove each component or substituted it with an alternative within the full model. This results in six distinct configurations for analysis: (1) Full Components: Our complete method with all components; (2) Random Selection: The method that randomly select conversion prompts from candidates instead of using our selection strategy (Sec. 3.3); (3) Class Prompt: The method categorizes the confusion scores of conversion prompts into three levels: high, medium, or low. These categorized labels are then used as in-context demonstrations for generating adversarial style-conversion prompts (Sec. 3.2), instead of relying on continuous confusion scores $s_{c}$ (Eq. 1); (4) Adversarial only Selection: The method that selects top-$k$ adversarial prompts, not considering diversity and coherence criteria; (5) w/o Adversarialness: The method that omits the adversarialness criterion in our selection strategy; (6) w/o Coherency: The method that omits the coherency criterion in our selection strategy; (7) w/o Score trajectory: The method without score trajectory component for the style-conversion prompt generation (Sec. 3.2).
@@
Figure 5. Performance of selection strategies over training rounds on PolitiFact. The x-axis represents the training rounds, while the y-axis represents the detector’s AUC.
Table 3 demonstrates that omitting any component leads to a decrease in performance for certain style conversions attack. Notably, excluding the score trajectory component when selecting the style-conversion prompt proved to be the most detrimental to performance. This finding suggests that the LLM can effectively identify conversion prompts that may confuse the detector through the style-conversion prompt and confusion score pairs as in-context demonstrations. Moreover, selectively choosing conversion prompts identified by the LLM based on specific criteria resulted in further performance improvements. Our sampling strategy also facilitates faster convergence compared to alternative sampling methods (See Figure 5).
Table 4. Performance comparison of different LLM-based baselines on the PolitiFact Dataset. (NY: The New York Times, TS: The Sun, NE: National Enquirer)
Model CNN NY TS NE Clean
GPT-3.5 0.5820 0.6242 0.6173 0.5274 0.7037
GPT-4 0.7133 0.6748 0.7025 0.6021 0.7823
4.3. Performance Analysis
We here conduct analysis on how AdStyle demonstrates robust and high performance across various scenarios and how it effectively enhances the detector’s performance.
Comparison with LLM-based baselines. AdStyle enhanced the detector’s performance by leveraging the reasoning abilities of advanced large language models like GPT-3.5. To determine whether the capabilities of an advanced large language model alone are sufficient for the fake news detection task, we compared AdStyle with other LLM-based baselines: zero-shot and in-context learning-based inference with GPT-3.5. In the case of the in-context learning baseline, one example each of fake news and real news was provided as in-context demonstrations. The example instruction prompt for the LLM-based baselines are as follows:
Table 4 presents the comparison results on the PolitiFact dataset. Using LLMs with prompting alone (Figure 6) makes it challenging to accurately determine the authenticity of news. Instead of ‘directly’ leveraging the LLM’s text generation capability for inference, it is more effective to use it as an augmentation tool to provide additional training signals, as demonstrated by our model.
Figure 6. Instruction prompt for LLM-based baselines. The ”news article” section contains the original news text.
Table 5. Comparison under attack scenarios with Gemini-Pro on the PolitiFact dataset.
Model CNN NY TS NE Average
Vanilla 0.7913 0.8039 0.8882 0.8397 0.8308
UDA 0.7216 0.7863 0.8657 0.8119 0.7964
RADAR 0.8023 0.7805 0.8764 0.8423 0.8254
ENDEF 0.7730 0.7537 0.8439 0.8058 0.7941
SheepDog 0.8487 0.8926 0.9174 0.8998 0.8896
Ours 0.8821 0.9120 0.9295 0.9241 0.9120
Robustness against a other LLM backbones. We assessed the robustness by examining its performance against style conversion attacks when the attacker utilizes a different LLM backbone, such as Gemini-Pro (Reid et al., 2024). As shown in Table 5, AdStyle consistently outperforms the baselines even with a different LLM backbone, suggesting that the proposed approach is not overly reliant on recognizing the specific style of content generated by the backbone.
Robustness against other possible attack scenarios. To validate the model’s robustness against various attacks, we have conducted comparison experiments with diverse attack scenarios to deceive the detector by altering textual style of inputs: (1) Adversarial prompt: Given a news article and its label, the prompt instructs the LLM to rewrite the article to evade detection as the given label. (2) Summarization prompt: A prompt instructing the LLM to summarize the news article without incorporating stylistic guidance. (3) In-Context prompt: A prompt providing an example of a recent, real CNN article, instructing the LLM to rewrite the given article in the same style. (4) Adversarial Paraphraser: The attack utilizes a paraphraser adversarially trained on the training dataset. The example prompt for each attack is described in Figures 12 to 14 in Appendix. Based on the results in Table 6, we confirm that our proposed model still demonstrates a performance improvement compared to other baselines against these attacks.